OpenStudy (anonymous):

multiply and simplify.√6(√3+5√2) please help

OpenStudy (valpey):

You can use the distributive property to multiply through by sqrt(6): \[\sqrt6(\sqrt3+5\sqrt2)=\sqrt6*\sqrt3+\sqrt6*5\sqrt2)\] Next you can aggregate under the roots: \[\sqrt{6*3}+5\sqrt{6*2}=\sqrt{18}+5\sqrt{12}\] Next factor out squares which divide 18 or 12 (9 divides 18 and 4 divides 12) \[=\sqrt{9*2}+5\sqrt{4*3}=3\sqrt{2}+5*2\sqrt{3}=3\sqrt{2}+10\sqrt{3}\]